Please help mi 2 solve dis 2x-5y=0-2
Answer by Alan3354(69443)

2x-5y=0-2
2x-5y=-2
--------
1 equation in 2 variables has an infinite # of solutions.
Pick any value for x and find y, or vice versa.
-----
No unique solution.